OpenStudy (anonymous):

describe the transformation required to obtain the graph of the given function from the basic trigonometric grapg y=cscx-9

OpenStudy (anonymous):

is that y = csc(x-9) or y = csc(x) -9 ??

OpenStudy (anonymous):

y=csc(x-9)

OpenStudy (cwrw238):

the graph is translated to the right along the x axis by 9 units